OpenStudy (anonymous):

Find the indicated roots and graph them in the complex plane. The fifth roots of i From the graph give the exact roots. (Express θ in radians.)

OpenStudy (anonymous):

exp(i (n2pi + pi) /5 ) n=0,1,2,3,4 = exp(i*pi/5), exp(i3*pi/5), exp(i5*pi/5), exp(i7*pi/5), exp(i9*pi/5) = exp(i*pi/5), exp(i*pi*3/5), exp(i*pi), exp(i*pi*7/5), exp(i*pi*9/5) These are points on the unit circle at angles of pi/5, 3pi/5, pi, 7pi/5, 9pi/5